# Parameter-Efficient Fine-Tuning of Large Language Models for Unit Test Generation: An Empirical Study

### v1

@@ V-B RQ2: PEFT training resource utilization effectiveness
One of the main advantages of using PEFT methods is that they require significantly less computing during training, mainly because of reduced backpropagation calculations. The fewer trainable parameters, the more efficient. Results in Figure 2 show a significant reduction of trainable parameters of PEFT methods compared to full fine-tuning. This is also essential to fit big models into GPU memory. For example, we could train the biggest 16 billion parameter model using a single NVIDIA A100 80GB GPU for LoRA, whereas we needed 6 x NVIDIA A100 40GB GPU for full fine-tuning.
CodeGen-350M-multiCodeGen2-1BStarCoder2-3BCodeGen2-3.7BCodeLlama-7BCodeGen2-7BStarCoder2-7BStarCoderBaseStarCoder2-15BCodeGen2-16B1.01.0-1.0- 1.00.50.5-0.5- 0.50.00.00.00.00.50.50.50.51.01.01.01.0Relative importanceLoRA(IA)3Prompt tuning
Figure 3: Stacked bar chart of symmetrically normalized CodeBLEU score increase and decrease, relative to the fraction of trained parameters for each model. Positive values mean an increased CodeBLEU score, while negative values indicate a decreased CodeBLEU score.
@@
Figure 3 shows the CodeBLEU score of each tuning method relative to the fraction of trainable parameters, symmetrically normalized between -1 and 1. As shown in Figure 3, prompt tuning has the highest efficiency among seven (i.e., CodeGen-350M-multi, CodeGen2-1B, StarCoder2-3B, CodeGen2-3.7B, StarCoder2-7B, StarCoder2-15B, and CodeGen2-16B) of the ten models and is the method that you get the ”most bang for your buck.” Moreover, prompt tuning has the most potential effect on models with large sizes, i.e., StarCoder2-15B and CodeGen2-16B. This can be explained by the increased capacity of larger models to understand better and leverage the additional context provided by prompt tuning. However, prompt tuning does not show consistent superiority for all the models. The method contributes to negative performance in three models (CodeLlama-7B, CodeGen2-7B, and StarCoderBase).
@@
For the three smallest models (CodeGen-350M-multi, CodeGen2-1B, and StarCoder2-3B), (IA)3 clocks in at second place, which provides better results than LoRA. However, (IA)3 fails to improve the three largest models. LoRA achieves a relatively low efficiency compared to the other PEFT methods. However, it does not lead to a negative impact on the performance of all models. Full fine-tuning obviously has the worst efficiency, as it needs to train all the model parameters.
@@
Summary for RQ2: PEFT methods offer significantly reduced computational requirements compared to full fine-tuning and are ranked by resource utilization effectiveness as follows: prompt tuning $>$ (IA)3 $>$ LoRA $>$ full fine-tuning. Prompt tuning demonstrates the highest returns on investment for most models but also carries risks of negative performance impact.
VI Discussion
@@
VI-A Comparison with related works
@@
Unlike existing studies, we investigated using PEFT and LLM to generate complete unit tests because the effectiveness of PEFT methods is task-dependent. Our results provide novel insights into their performance in the context of unit test generation. For example, Ayupov and Chirkova [5] conclude that LoRA is less effective than full fine-tuning for code generation. We show contradictory results that LoRA is more effective than fine-tuning in unit test generation. Weyssow et al. [8] explored generating code from natural language instructions and concluded that LoRA is the most effective one in the PEFT methods they evaluate. Our results show that prompt tuning is the most effective in terms of cost and resource utilization in unit test generation, although LoRA is more effective without considering the cost aspects.
Compared to the existing studies, we also performed a more comprehensive evaluation, as we invested significant efforts to compare all PEFT methods to full fine-tuning across all model sizes (up to 16B parameters). In contrast, Ding et al. [24] only evaluated PEFT using the 355 million parameter RoBERTaLARGE model, while Weyssow et al. [8] only did full fine-tuning for smaller models up to maximum 350 million parameters.
@@
Several studies have leveraged LLMs for generating unit tests [1, 2, 3, 4, 29]. However, none of the works investigate the potential performance of using PEFT compared to full fine-tuning. Tufano et al. [1] automatically generated unit tests and reported 95% syntactic correctness when training BART large (400 million parameters) [52] on Methods2Test. With PEFT methods, we can generate an average median of 94.03% syntactically correct unit tests. Training between 0.02 million and 13 million parameters on 1% of the same data showcasing the potential powerful performance of PEFT methods. Our results also illustrate that LoRA surpasses full fine-tuning in terms of boosting the quality of the generated unit tests measured using CodeBLU in several cases.
VI-B Implications to Academia and Industry
@@
Writing unit tests is a hard and often time-consuming task [53]. However, they are essential to ensure code quality, stability, and maintainability [54]. Automating unit test generation using LLMs can significantly reduce the time and effort required for this process, making it an attractive solution for academia and industry. However, as our experiments show, full fine-tuning is sometimes required to achieve adequate quality of the generated unit tests, which is computationally demanding. Many LLM users may not have the necessary computing resources to fine-tune LLMs fully. Our results show that various PEFT methods can achieve similar performance as full fine-tuning by training a minuscule parameter fraction. Furthermore, our results suggest starting with prompt tuning for cost-effectiveness. Then, if stability issues arise, LoRA should be considered as a reliable alternative because it does not lead to negative performance impact.
VI-C Threaths to Validity
VI-C1 External validity
@@
One of the threats to external validity is the selection of models. We minimize this threat by selecting a broad range of models varying in size, pre-training data, and learning objectives. Another potential concern is the validity of the Methods2Testsmall benchmarking dataset. We opted for this dataset over other alternatives because it comprises real-world unit tests. We also validate the Java-translated version of the popularly used HumanEval code benchmarking dataset to mitigate the threat. The last threat is related to our evaluation metric. While popularly used, we recognize BLEU as a sub-optimal performance estimator for code generation. We, therefore, employ a more code-friendly version, CodeBLEU. Although CodeBLEU may not be the optimal metric to evaluate the functional correctness of unit tests completely, it provides a good tradeoff between evaluation effort and validity. To supplement our evaluation, we also calculate pass@1 and code coverage for the Methods2Testsmall dataset, as this contains isolated code samples that are easier to test.
VI-C2 Internal validity
@@
The main threat to internal validity is comprised of the selection of hyperparameters. We use the hyperparameters reported in the introductory paper of each PEFT method. Additionally, we use greedy decoding to support reproducibility and reduce variability in the generated results.
VII Conclusion and future work
@@
Although PEFT has been investigated to boost the quality and efficiency of code-related tasks using LLMs, they have not been studied in the context of unit test generation. We performed the first empirical study to compare different PEFT methods and full fine-tuning to generate unit tests. Our comparison involved ten LLMs of different sizes and architectures. Our results offer valuable insights into the advantages and challenges of using PEFT and full fine-tuning for unit test generation. With the knowledge we’ve gathered, individuals can choose the most suitable PEFT methods or full fine-tuning based on the size and characteristics of the LLMs.
@@
VIII Data availability
The Methods2Testsmall dataset is available at:
https://huggingface.co/datasets/andstor/methods2test_small
All fine-tuned models are also made available at:
https://huggingface.co/andstor/peft-unit-test-generation-experiments
Experiment code and results are available at:
https://github.com/andstor/peft-unit-test-generation-replication-package

### v2

@@ 5.2 RQ2: PEFT training resource utilization effectiveness
One of the main advantages of using PEFT methods is that they require significantly less computing during training, mainly because of reduced backpropagation calculations. The fewer trainable parameters, the more efficient. As seen from the second column in Section 5.1(a), PEFT methods show a significant reduction of trainable parameters compared to full fine-tuning. This is also essential to fit big models into GPU memory. For example, we could train the largest 16 billion parameter model using a single NVIDIA A100 80GB GPU for LoRA, whereas we needed 6 x NVIDIA A100 40GB GPUs for full fine-tuning.
Figure 2: Stacked bar chart of symmetrically normalized metrics value differences (increase or decrease) between baseline and tuned models, divided by the percentage of trainable parameters of models, on the Methods2Testrunnable dataset.
@@
Figure 2 shows the results of the metrics (valid syntax, CodeBLEU, pass@1, instruction coverage, branch coverage, and mutation score) for each tuning method relative to the percentage of trainable parameters, symmetrically normalized between -1 and 1. As shown in Figure 2, prompt tuning has the highest efficiency for every metric and is the method that you get the “most bang for your buck.” Moreover, prompt tuning has the most potential cost-effectiveness on models with large sizes, i.e., Qwen2.5-Coder-14B, StarCoderBase, StarCoder2-15B, and CodeGen2-16B. This can be explained by the increased capacity of larger models to better understand and leverage the additional context provided by prompt tuning. However, prompt tuning does not show consistent superiority for all the models. The method contributes to negative performance in three models (StarCoder2-3B, CodeLlama-7B, and CodeGen2-7B).
@@
For the four smallest models (Qwen2.5-Coder-3B, CodeGen-350M-multi, CodeGen2-1B, and StarCoder2-3B), (IA)3 clocks in at second place, which provides better results than LoRA. However, (IA)3 generally fails to efficiently improve the three largest models. LoRA achieves a relatively low efficiency compared to the other PEFT methods. However, with the exception of pass@1, it does not lead to any significant negative impact on the performance of any of the models. Full fine-tuning obviously has the worst efficiency, as it needs to train all the model parameters.
@@
Summary for RQ2: PEFT methods offer significantly reduced computational requirements compared to full fine-tuning and are ranked by resource utilization effectiveness as follows: prompt tuning >> (IA)3 >> LoRA >> full fine-tuning. Prompt tuning demonstrates the highest returns on investment for most models but also carries risks of negative performance impact.
6 Discussion
@@
6.1 Comparison with related studies
@@
As the effectiveness of PEFT methods is task-dependent, our results provide novel insights into their performance in the context of unit test generation. [Ayupov2022ParameterEfficientFO] conclude that LoRA is less effective than full fine-tuning for code generation. We show contradictory results that LoRA could be as good as or even more effective than fine-tuning in unit test generation. [weyssow2024exploring] explored generating code from natural language instructions and concluded that LoRA is the most effective among the PEFT methods they evaluated. Our results show that prompt tuning is the most effective in terms of cost and resource utilization in unit test generation, although LoRA provides better performance without considering the cost aspects.
Compared to the existing studies, we also performed a more comprehensive evaluation, as we invested significant efforts to compare all PEFT methods to full fine-tuning across all model sizes (up to 16 billion parameters). In contrast, [Ding2023ParameterefficientFO] only evaluated PEFT using the 355 million parameter RoBERTaLARGE model, while [weyssow2024exploring] only did full fine-tuning for smaller models up to a maximum of 350 million parameters. While existing works on code generation or test generation have demonstrated promising results, they are often limited to artificially constructed benchmarks such as HumanEval. These datasets are typically small, synthetically curated, and lack the complexity of real-world software systems. As our results show, such controlled settings also overlook critical challenges that arise in real-world scenarios, such as missing dependencies, inconsistent, or erroneous code. Defects4J [just2014defects4j], one of the more realistic benchmarks available, contains only 17 repositories. This starkly contrasts the scale and messiness of actual industrial or open-source projects. As a result, prior work such as [alagarsamy2024a3test] may significantly overestimate the robustness and generalizability of models when deployed beyond synthetic testbeds. Our evaluation, grounded in more realistic settings using Methods2Test [tufano2021unit], highlights these limitations and emphasizes the need for evaluation frameworks that reflect the true complexity of software development.
@@
Several studies have leveraged LLMs for generating unit tests [tufano2021unit, siddiq2024using, yuan2023manual, schäfer2023empirical, steenhoek2023reinforcement]. However, none of the works investigate the potential performance of using PEFT compared to full fine-tuning. [tufano2021unit] automatically generated unit tests and reported 95% syntactic correctness when training BART large (400 million parameters) [lewis2019bart] on Methods2Test. With PEFT methods, we can generate an average median of 96% syntactically correct unit tests. Training between 20 thousand and 13 million parameters on 1% of the same data, showcasing the potentially powerful performance of PEFT methods.
Different from studies [siddiq2024using, yuan2023manual, schäfer2023empirical] using large instruction-tuned enterprise LLMs such as ChatGPT to generate test cases, this study focused on generating test cases from autoregressive base code models. Compared to e.g. [schäfer2023empirical], our best model produced 13% less statement coverage and 55% lower branch coverage. Achieving state-of-the-art performance was not the focus of our work. Instead, our work provides an empirical evaluation of PEFT methods in a more practical scenario, where organizations can efficiently customize models for their own testing needs. One reason for the reduced coverage is probably due to the size differences of the models. The largest model we evaluated had 16 billion paramaters, opposed to the lager commercial GPT-3.5-Turbo. Our results shown in Section 5.1 illustrate that larger models generally generate better unit tests than smaller ones. Another reason is probabaly due to the differences between the benchmark datasets used by our study. We use a real-world evaluation dataset that is more diverse, spanning 119 repositories, versus the 25 npm packages used by [siddiq2024using]. Our analyses on the reasons for low pass@1 rates show that it is necessary to consider potential compilation and code dependency issues when using LLMs to generate executable unit tests, providing valuable novel insights to improve unit test generation using PEFT or ChatGPT. Opposed to related works like [siddiq2024using, yuan2023manual, schäfer2023empirical], we conduct mutation testing. As our results depict in section 5.1.5, measuring mutation score is crucial to ensure generated tests are actually able to reveal faults.
6.2 Implications to Academia and Industry
@@
Writing unit tests is a hard and often time-consuming task [7102609]. However, they are essential to ensure code quality, stability, and maintainability [Gonzalez2017ALS]. Automating unit test generation using LLMs can significantly reduce the time and effort required for this process, making it an attractive solution for academia and industry. As our experiments show, full fine-tuning is sometimes required to achieve adequate quality of the generated unit tests, which is computationally demanding. Many LLM users may not have the necessary computing resources to fine-tune LLMs fully. Our results show that various PEFT methods can achieve similar performance to full fine-tuning by training a minuscule parameter fraction. Furthermore, our results suggest starting with prompt tuning for cost-effectiveness. Then, if stability issues arise, LoRA should be considered as a reliable alternative because it does not lead to a negative performance impact.
6.3 Threats to Validity
6.3.1 External validity
@@
One of the threats to external validity is the selection of models. We minimize this threat by selecting a broad range of models varying in size, pre-training data, and learning objectives. As we have have spent upward of 700 GPU hours for model tuning alone, exploring an even broader range of alternative models is constrained due to cost. Another potential concern is the validity of the benchmarking dataset. We chose to use Methods2Test, which is the largest publicly available corpus of real-world Java unit test cases and corresponding focal methods. We believe this provides a representative evaluation setting compared to existing studies like [alagarsamy2024a3test], although future work should expand the analysis to other programming languages to ensure generalizability. The last threat is related to our evaluation metric. While popularly used, we recognize BLEU as a sub-optimal performance estimator for code generation. We, therefore, employ a more code-friendly version, CodeBLEU. Although CodeBLEU may not be the optimal metric to evaluate the functional correctness of unit tests completely, it provides a good tradeoff between evaluation effort and validity. To supplement our evaluation, we have also spent significant effort developing evaluation pipelines, as well as over 200 CPU hours for executing the test in the Methods2Testrunnable dataset, calculating pass@1, code coverage, and mutation scores.
6.3.2 Internal validity
@@
The main threat to internal validity is comprised of the selection of hyperparameters. We use the hyperparameters reported in the introductory paper of each PEFT method to avoid introducing unnecessary biases. Additionally, we use greedy decoding to support reproducibility and reduce variability in the generated results.
7 Conclusion and Future Work
@@
Although PEFT has been investigated to boost the quality and efficiency of code-related tasks using LLMs, they have not been studied in the context of unit test generation. We performed the first empirical study to compare different PEFT methods and full fine-tuning to generate unit tests. Our comparison involved thirteen LLMs of different sizes and architectures. Our results offer valuable insights into the advantages and challenges of using PEFT and full fine-tuning for unit test generation. With the knowledge we’ve gathered, individuals can choose the most suitable PEFT methods or full fine-tuning based on the size and characteristics of the LLMs. In future work, we plan to extend the analysis perform an empirical comparison of methods for tuning instruction-based LLMs, covering both PEFT methods per se and variations in prompt templates
@@
Acknowledgements
The empirical results were supported through the computational resources of HPC facilities at the Norwegian University of Science and Technology (NTNU) [sjalander2019epic].
Data Availability
The Methods2Testsmall dataset is available at: https://huggingface.co/datasets/andstor/methods2test_small
The Methods2Testrunnable dataset is available at: https://huggingface.co/datasets/andstor/methods2test_runnable
The Methods2Testmeta dataset is available at: https://huggingface.co/datasets/andstor/methods2test_meta
Metadata and links to the trained models can be found at: https://huggingface.co/datasets/andstor/peft-unit-test-generation-experiments
Experiment code and results are available at: https://github.com/andstor/peft-unit-test-generation-replication-package